OpenStudy (turingtest):

How many numbers end in the four digits 1995 and become an integer number of times smaller when these digits are erased?

OpenStudy (turingtest):

I guess we at least need the factors of 1995 3*5*7*19 that's a start...

OpenStudy (anonymous):

of the form 1995n/(n-1)where n is integer...and the no ends in 1995..trial

OpenStudy (turingtest):

It is as I posted it It seems to mean something like how many numbers when changed from \[abcd1995 \to abcd\] decrease by a factor that is an integer.

OpenStudy (turingtest):

No, an old magazine called "quantum" has these. They don't sell that magazine anymore :(

OpenStudy (turingtest):

it is sitting in the meta-math section in case nobody has the time to solve it here...

OpenStudy (turingtest):

not as far as I can see

OpenStudy (zarkon):

like this \(19951995\to1995\)

OpenStudy (turingtest):

so it is! 10001 how did zarkon know?

OpenStudy (zarkon):

10001,100001,.......

OpenStudy (zarkon):

an educated guess

OpenStudy (turingtest):

Oh I think I see where this is going...

OpenStudy (turingtest):

but that sequence is infinite

OpenStudy (zarkon):

yes

OpenStudy (turingtest):

I'm pretty sure the answer should be finite just by tendency of this magazine not to have trick questions like that.

OpenStudy (turingtest):

what's that?

OpenStudy (turingtest):

oh yeah, heard of it but never tried it I'll be patient and let it sit in meta-math though

OpenStudy (zarkon):

so we have one so far :) I believe I now understand the full problem

OpenStudy (turingtest):

I had a thought and lost it... I'd appreciate it if you could walk me through your thought process zarkon, if you come up with something.

OpenStudy (anonymous):

abcd1995 --> abcd Let x be the number abcd. Let y be the integer number of times 10000x + 1995 = abcd1995 10000x + 1995 = xy y = (10000x + 1995)/x (10000x + 1995) must be divisible by x

OpenStudy (mathmate):

Would 19951995 be an answer, since 19951995/1995=10001

OpenStudy (anonymous):

x must be one of the factors of 1995

OpenStudy (anonymous):